# MERGE: Matching Electronic Results with Genuine Evidence for verifiable voting in person at remote locations

### v3

@@ B.6 Zero-knowledge proofs
Principals: The prover $P$ and the verifier $V$
Public input: ElGamal parameters $\mathcal{G}=(p,q,g)$, ciphertext $(\alpha,\beta)=(g^{\xi},KK^{\xi})$, public key $K$, two deployment-dependant constants $cons_{0}$ and $cons_{1}$
Private input ($P$): $\xi$ s.t. $(\alpha,\beta)=(g^{\xi},KK^{\xi})$
  1. 1.
    $P$:
    • $u_{0},u_{1},c_{0}\xleftarrow{\$}\mathbb{Z}_{q}$
    • $(a_{0},b_{0})=(g^{u_{0}}\mod p,K^{u_{0}+c_{0}}\mod p)$, $(a_{1},b_{1})=(g^{u_{1}}\mod p,K^{u_{1}}\mod p)$
    • $c=H(cons_{0};cons_{1},K,\alpha,\beta,a_{0},b_{0},a_{1},b_{1})$
    • $c_{1}=c-c_{0}$, $v_{0}=u_{0}-c_{0}\xi\mod q$, $v_{1}=u_{1}-c_{1}\xi\mod q$
    • Send $(c_{0},c_{1},v_{0},v_{1})$ to $V$
  2. 2.
    $V$
    1. (a)
      $(a_{0},b_{0}):=(g^{v_{0}}\alpha^{c_{0}}\mod p,K^{v_{0}}\beta^{c_{0}}\mod p)$
    2. (b)
      $(a_{1},b_{1}):=(g^{v_{1}}\alpha^{c_{1}}\mod p,K^{v_{1}-c_{1}}\beta^{c_{1}}\mod p)$
    3. (c)
      Verify if $c_{0}+c_{1}=H(cons_{0};cons_{1},K,\alpha,\beta,a_{0},b_{0},a_{1},b_{1})$. Otherwise, reject.
Required operations: 4 exp
Required operation for verification: 4 multi-exp
B.7 Verification Summary
@@
B.7.1 Cast-as-intended verification
@@
B.7.2 BB transcript verification (public)
@@
    Verify that no two ballots on the bulletin board share the same ciphertext (only necessary for privacy. See Section 5.2).
@@
B.7.3 Verification at the Local Counting Center
@@ B.8 Failure handling
B.8.1 Cast-as-intended verification
@@
B.8.2 BB transcript verification (public)
@@
Item 2: Given the negligible probability that any two ciphertexts are identical, if a ciphertext matches an already published one, it is highly likely due to the voting machine’s fault.
@@
B.8.3 Verification at the Local Counting Center
@@ C.1 Verifiability
Differences between this world and the Intermediate World 5 are as follows: The serial numbers and the vote contents of digital ballots are initially recorded by the voting machine in an encrypted format. Each digital ballot is also accompanied by a zero-knowledge proof demonstrating the validity of the (encrypted) vote. Subsequently, all digital ballots undergo the same processing steps as those within MERGE (i.e., steps 1 to 5 from the digital path explained in Section 4.2) followed by private decryption in step 9.
@@
    • decryption of the encrypted tally obtained at step 3 of the digital path (i.e., Section B.7.2, BB transcript verification (public), item 7)
@@
  • aggregation of encrypted ballots at step 2 of the digital path (i.e., Section B.7.2, BB transcript verification (public), item 6)
@@
Let $\mathcal{B}$ denote the list of digital ballots recorded on the BB for $N$ voters before initiation of the step 1 in the Intermediate World 6:
@@
be the digital ballots recorded by the voting machine on the BB in Intermediate World 5. The tally in Intermediate World 5 ($\sum_{i=1}^{N}\mathbf{b_{i}}$) matches the result of decryption at step 2 of the protocol in the Intermediate World 6 ($\mathit{D}(\prod_{j=1}^{N}\mathit{E}(\mathbf{b^{\prime}_{j}}))$). This equivalence is due to the homomorphic properties of the encryption algorithm. The private decryption at step 3 of the protocol in the Intermediate World 6 results in $\sum_{j=1}^{N}\mathbf{b^{\prime}_{j}}$. Otherwise the security of the proof of decryption protocol would be violated. Since the paper path in both worlds is the same, the paper manifest in Intermediate World 5 can be selected to match that in Intermediate World 6. Each sequence of ballots in Intermediate World 6 corresponds to a sequence in Intermediate World 5 (established through the $\pi^{-1}$ permutation function), maintaining identical discrepancy values. Consistency in both discrepancy values and tally across these intermediate worlds ensures an equivalent RLA risk. ∎
@@
Let’s start with the setup in Intermediate World 3. In this world, each digital ballot is categorized as either ‘not matching’, ‘one-to-one’, or ‘duplicated’. For each digital ballot in the ‘duplicated’ category, one of the paper ballots with a matching serial number that results in the lowest discrepancy is identified as the corresponding paper ballot for that digital ballot. Let the set of digital ballots and those in either the ‘one-to-one’ or ‘duplicated’ category be denoted by $\mathcal{B}$ and $\mathcal{B}^{\prime}$, respectively. Let the set of paper ballots and those corresponding to $\mathcal{B}^{\prime}$ be denoted by $\mathcal{P}$ and $\mathcal{P}^{\prime}$, respectively. The discrepancy for each ballot in $\superscript\mathcal{B}\backslash\mathcal{B}^{\prime}$ is set to the maximum possible value for that ballot.
@@ C.2.1 Assumptions
The first assumption is that the basic ElectionGuard protocol satisfies BPRIV.
@@ C.2.2 The proof
Valid$(BB_{\beta},b_{\beta})$ Verifies the vote validity ZKPs and digital signatures for each vote, then removes duplicates (as specified in Digital Processing Step 1).
@@
$\mathsf{Simproof(BB_{1},r)}$ takes the true aggregated vote ciphertext from $BB_{1}$, the true mixing proof (input ciphertexts, output ciphertexts and $\textit{ZKP}_{\textsc{Mix}}(,)$) from $BB_{1}$, and the decrypted votes from $r$, and uses the ZK simulator of the decryption ZKP to simulate both the decryptions of the aggregate (from Step 3) and the individual vote decryptions (from Step 5).
@@
First recall the definition of NM-CPA
[…]
NM-CPA is defined by the following game:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    The adversary submits a guess $\beta^{\prime}$ for $\beta$.
@@
Consider what the adversary could have used for $b_{i}$ on board $BB_{1}$. If it is identical to $b_{i}$ from $BB_{0}$ then distinguishing is impossible. If it is different, then either the adversary re-uses a ciphertext already on the board (in which case, the removal of duplicate ciphertexts corresponds to the return of $\bot$ in the NM-CPA game, no information is returned, and distinguishing is impossible), or the game is run and distinguishing the two boards—which differ only in one ballot—is equivalent to distinguishing those two ballots in the NM-CPA game. None of these options give $\mathcal{A}$ a better advantage than distinguishing the two ballots in the NM-CPA game, which we assume to be negligible.
@@ D.1 VAULT-MERGE: incorporating MERGE in to an RLA that uses VAULT for all ballots
  • if the ciphertext is a MERGE ballot, it should be treated as described in subsubsection 4.4.1, using a ballot paper of matching serial number if possible, and a worst-case assumption if there is no ballot paper,
@@ D.2 Sub-MERGE: hiding MERGE ballots in larger batches
Call the (small) set of MERGE ballots on the BB $M$. When $|M|$ is small, mixing does not have much benefit. We could use the ballots output from the mix, but it is just as valid to skip mixing and use the aggregate of $M$ from before the mixing step that is published in Step 2 of the digital path.

### v4

@@ B.2 Notations and building blocks
See [6] for details of these zero knowledge proofs. These are applied for our “ZKP of proper ballot construction”.
B.3 Verification Summary
@@
B.3.1 Cast-as-intended verification
@@
B.3.2 BB transcript verification (public)
@@
    Verify that no two ballots on the bulletin board contain a common ciphertext (only necessary for privacy. See Section 5.2).
@@
B.3.3 Verification at the Local Counting Center
@@ B.4 Failure handling
B.4.1 Cast-as-intended verification
@@
B.4.2 BB transcript verification (public)
@@
Item 2: Given the negligible probability that any two ciphertexts are identical, if a ciphertext matches an already published one, it is overwhelmingly likely to be the voting machine’s fault (though if the duplicated ciphertexts come from two different voting computers, it is not clear which one is at fault, unless there is reliable timing information).
@@
B.4.3 Verification at the Local Counting Center
@@ C.1 Verifiability
Differences between this world and the Intermediate World 5 are as follows: The serial numbers and the vote contents of digital ballots are initially recorded by the voting machine in an encrypted format. Each digital ballot is also accompanied by a zero-knowledge proof demonstrating the validity of the (encrypted) vote. Subsequently, all digital ballots undergo the same processing steps as those within MERGE (i.e., steps 10 to 6 from the digital path explained in Section 4.2) followed by private decryption in step 9.
@@
    • decryption of the encrypted tally obtained at step 4 of the digital path (i.e., Section B.3.2, BB transcript verification (public), item 7)
@@
  • aggregation of encrypted ballots at step 3 of the digital path (i.e., Section B.3.2, BB transcript verification (public), item 6)
@@
Let $\mathcal{B}$ denote the list of digital ballots recorded on the BB for $N$ voters before initiation of the step 10 in the Intermediate World 6:
@@
be the digital ballots recorded by the voting machine on the BB in Intermediate World 5. The tally in Intermediate World 5 ($\sum_{i=1}^{N}\mathbf{b_{i}}$) matches the result of decryption at step 3 of the protocol in the Intermediate World 6 ($\mathit{D}(\prod_{j=1}^{N}\mathit{E}(\mathbf{b^{\prime}_{j}}))$). This equivalence is due to the homomorphic properties of the encryption algorithm. The private decryption at step 4 of the protocol in the Intermediate World 6 results in $\sum_{j=1}^{N}\mathbf{b^{\prime}_{j}}$. Otherwise the security of the proof of decryption protocol would be violated. Since the paper path in both worlds is the same, the paper manifest in Intermediate World 5 can be selected to match that in Intermediate World 6. Each sequence of ballots in Intermediate World 6 corresponds to a sequence in Intermediate World 5 (established through the $\pi^{-1}$ permutation function), maintaining identical discrepancy values. Consistency in both discrepancy values and tally across these intermediate worlds ensures an equivalent RLA risk. ∎
@@
Let’s start with the setup in Intermediate World 3. In this world, each digital ballot is categorized as either ‘not matching’, ‘one-to-one’, or ‘duplicated’. We will prove that, in each of these categories, each ballot’s discrepancy is maintained or decreased in the transition to Intermediate World 2. Figure 5 shows an example of the transition.
For each digital ballot in the ‘duplicated’ category, one of the paper ballots with a matching serial number that results in the minimum discrepancy is identified as the corresponding paper ballot for that digital ballot, and this correspondence is retained in Intermediate World 2. Remember that “minimum discrepancy” means “the minimum discrepancy number, where 0 is taken to be a match, overstatements are positive, and understatements are negative.” Because we were choosing the “minimum discrepancy” for our RLA accounting in Intermediate World 3, it’s obvious that these one-to-one matches we just made in Intermediate World 2 produces exactly the same discrepancies for every digital ballot in the ‘duplicated’ category.
[…]
For digital ballots in the ‘one-to-one’ category in Intermediate World 3, there is no change in Intermediate World 2. In Figure 5, Serial numbers 05, 11 and 15 are in this category.
Finally, consider how to deal with digital ballots that are in the ‘not matching’ category in Intermediate World 3. We want to make a one-to-one correspondence with the ballot papers, of which we must (by assumption) have no more than we have electronic records. So we just make one up! We pick an arbitrary assignment of as-yet-not-matched digital records to as-yet-not-matched papers. Paper serial numbers are edited if the paper’s serial number matched no digital record, or if it was a duplicate that was not selected as the minimum when dealing with the ‘duplicated’ category. Papers are added to make the number of digital and paper records equal, each with a serial number to match an as-yet-unmatched digital ballot. [^17] This is just a proof technique to demonstrate that the treatment of ballots is conservative. There is no actual alteration of serial numbers on ballots, nor addition of new ballot papers during the protocol.
In Figure 5, this arbitrary assignment is performed for $\textit{SN}=12$ (which is matched to the unused duplicate of $\textit{SN}=02$), for $\textit{SN}=18$ (which is matched to the previously-unmatched $\textit{SN}=31$) and for $\textit{SN}=21$ (which is matched to an invented ballot with an arbitrary vote selection). In each case, the serial number of the paper record is updated to match is corresponding digital one.
We need to argue that the RLA discrepancies must be no greater for these ballots, but that’s easy because when we had no match (in Intermediate World 3) we assumed the worst-case discrepancy. Now we have a specific match for each of them, which can’t, by definition, be worse than the worst case. (edited)
So, overall, this process produces a one-to-one match between digital and paper ballots, in which no digital record’s discrepancy has increased from Intermediate World 3 to Intermediate World 2.
be denoted by $\mathcal{B}$ and $\mathcal{B}^{\prime}$, respectively. Let the set of paper ballots and those corresponding to $\mathcal{B}^{\prime}$ be denoted by $\mathcal{P}$ and $\mathcal{P}^{\prime}$, respectively. The discrepancy for each ballot in $\superscript\mathcal{B}\backslash\mathcal{B}^{\prime}$ is set to the maximum possible value for that ballot.
@@ C.2.1 Assumptions
The first assumption is that the basic ElectionGuard encryption scheme, including its validity ZKPs, is non-malleable. This applies to the encryption of an individual ciphertext. We first recall the definition of non-malleability (NM-CPA).
[…]
NM-CPA is defined by the following game:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    The adversary submits a guess $\beta^{\prime}$ for $\beta$.
@@ C.2.2 The proof
Valid$(BB_{\beta},b_{\beta})$ Verifies the vote validity ZKPs and digital signatures for each vote, then removes any vote that shares a ciphertext with any other (as specified in Digital Processing Step 10).
@@
$\mathsf{Simproof(BB_{1},r)}$ takes the true aggregated vote ciphertext from $BB_{1}$, the true mixing proof (input ciphertexts, output ciphertexts and $\textit{ZKP}_{\textsc{Mix}}(,)$) from $BB_{1}$, and the decrypted votes from $r$, and uses the ZK simulator of the decryption ZKP to simulate both the decryptions of the aggregate (from Step 4) and the individual vote decryptions (from Step 6).
@@
Consider what the adversary could have used for $b_{i}$ on board $BB_{1}$. If it is identical to $b_{i}$ from $BB_{0}$ then distinguishing is impossible. If it is different, then either the adversary re-uses a ciphertext already on the board (in which case, the removal of duplicate ciphertexts corresponds to the return of $\bot$ in the NM-CPA game, no information is returned, and distinguishing is impossible), or the game is run and distinguishing the two boards—which differ only in one ballot—is equivalent to distinguishing those two ballots in the NM-CPA game. None of these options give $\mathcal{A}$ a better advantage than distinguishing the two ballots in the NM-CPA game, which we assume to be negligible (Assumption C.2.1).
@@ D.1 VAULT-MERGE: incorporating MERGE in to an RLA that uses VAULT for all ballots
  • if the ciphertext is a MERGE ballot, it should be treated as described in subsection 4.4, using a ballot paper of matching serial number if possible, and a worst-case assumption if there is no ballot paper,
@@ D.2 Sub-MERGE: hiding MERGE ballots in larger batches
Call the (small) set of MERGE ballots on the BB $M$. When $|M|$ is small, mixing does not have much benefit. We could use the ballots output from the mix, but it is just as valid to skip mixing and use the aggregate of $M$ from before the mixing step that is published in Step 3 of the digital path.